Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

A florist has 2 azaleas, 3 buttercups, and 4 petunias. She puts two flowers together at random in a bouquet. However,

Expert replies
by VJesus12 » Thu Sep 03, 2020 6:29 am
Timer
00:00
Your answer
Difficulty–

right wrong

Pick an answer to see how members did

A florist has 2 azaleas, 3 buttercups, and 4 petunias. She puts two flowers together at random in a bouquet. However, the customer calls and says that she does not want two of the same flower. What is the probability that the florist does not have to change the bouquet?

A. 5/18
B. 13/18
C. 1/9
D. 1/6
E. 2/9

Answer: B

Source: Manhattan GMAT
Join the discussion
Source: — Problem Solving |

VJesus12 wrote: ↑
Thu Sep 03, 2020 6:29 am
A florist has 2 azaleas, 3 buttercups, and 4 petunias. She puts two flowers together at random in a bouquet. However, the customer calls and says that she does not want two of the same flower. What is the probability that the florist does not have to change the bouquet?

A. 5/18
B. 13/18
C. 1/9
D. 1/6
E. 2/9

Answer: B

Source: Manhattan GMAT
First, we can rewrite the question as "What is the probability that the two flowers are different colors?"

Well, P(different colors) = 1 - P(same color)

Aside: let A = azalea, let B = buttercup, let P = petunia

P(same color) = P(both A's OR both B's OR both P's)
= P(both A's) + P(both B's) + P(both P's)

Now let's examine each probability:

P(both A's):
We need the 1st flower to be an azalea and the 2nd flower to be an azalea
So, P(both A's) = (2/9)(1/8) = 2/72

P(both B's)
= (3/9)(2/8) = 6/72

P(both P's)
= (4/9)(3/8) = 12/72

So, P(same color) = (2/72) + (6/72) + (12/72) = 20/72 = 5/18

Now back to the beginning:
P(diff colors) = 1 - P(same color)
= 1 - 5/18
= 13/18

Answer: B

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

VJesus12 wrote: ↑
Thu Sep 03, 2020 6:29 am
A florist has 2 azaleas, 3 buttercups, and 4 petunias. She puts two flowers together at random in a bouquet. However, the customer calls and says that she does not want two of the same flower. What is the probability that the florist does not have to change the bouquet?

A. 5/18
B. 13/18
C. 1/9
D. 1/6
E. 2/9

Answer: B

Source: Manhattan GMAT
Total flowers: \(AA, BBB, PPPP\)

Probability that two same flowers are picked:

\(AA = \dfrac{2}{9} \cdot \dfrac{1}{8} = \dfrac{1}{36}\)

\(BB = \dfrac{3}{9} \cdot \dfrac{2}{8} = \dfrac{1}{12}\)

\(PP = \dfrac{4}{9} \cdot \dfrac{3}{8} = \dfrac{1}{6}\)

Then,
\(\dfrac{1}{36} + \dfrac{1}{12} + \dfrac{1}{6} = \dfrac{5}{18}\)

And
\(1 - \dfrac{5}{18} = \dfrac{13}{18}\)

Therefore, B
Join the discussion

VJesus12 wrote: ↑
Thu Sep 03, 2020 6:29 am
A florist has 2 azaleas, 3 buttercups, and 4 petunias. She puts two flowers together at random in a bouquet. However, the customer calls and says that she does not want two of the same flower. What is the probability that the florist does not have to change the bouquet?

A. 5/18
B. 13/18
C. 1/9
D. 1/6
E. 2/9

Answer: B

Solution:

Calculating the probability that the florist does not have to change the bouquet is the same as calculating the probability that the bouquet has two different flowers. We can use the following formula:

1 = P(getting two of the same flower) + P(not getting two of the same flower)

Let’s determine the probability of getting two of the same flower.

P(2 azaleas) = 2/9 x 1/8 = 2/72

P(2 buttercups) = 3/9 x 2/8 = 6/72

P(2 petunias) = 4/9 x 3/8 = 12/72

The probability of getting two of the same flower is, therefore, 2/72 + 6/72 + 12/72 = 20/72 = 5/18.

Thus, the probability of not getting two of the same flower is 1 - 5/18 = 13/18.

Alternate Solution:

Without any restrictions, there are 9C2 = 9!/(2!*7!) = (9*8)/2 = 36 ways to pick two flowers from a total of 9 flowers.

The two azaleas can be chosen in 2C2 = 1 way, two buttercups can be chosen in 3C2 = 3 ways and two petunias can be chosen in 4C2 = 4!/(2!*2!) = (4*3)/2 = 6 ways. In total, there are 1 + 3 + 6 = 10 ways to pick two flowers of the same kind and thus, 36 - 10 = 26 ways to pick two bouquets that contain two different kinds of flowers. We see that the probability that the selection of two flowers does not contain the same kind of flowers is 26/36 = 13/18.

Answer: B

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion